what is the chance of a boy being picked out if there are 360 girls and 350 boys if it is at random

Found 2 solutions by stanbon, oberobic:
Answer by stanbon(75887)

what is the chance of a boy being picked out if there are 360 girls and 350 boys if it is at random
---
# of ways to select a boy: 350
# of possible selections: 350+360 = 710
-----
P(select a boy) = 350/710 = 35/71
===================================
Cheers,
Stan H.

Answer by oberobic(2304)

There are 360 girls + 350 boys = 710 possibilities.
The are 350 favorable outcome, which means you pick a boy.
P(boy) = 350/710 = 35/71 = approximately .4929577464788732